# ADBM: Adversarial Diffusion Bridge Model for Reliable Adversarial Purification

### v1

@@ C.2 Additional Adaptive Attack Configurations
Appendix D Additional Evaluation Results
D.1 Comparison between Different Attack Methods on CIFAR-10
Similar to Lee and Kim [19], we first performed a comparison between PGD with 20 EOT samples (full gradient) and AutoAttack with the same number of EOT samples (the default setting of rand version AutoAttack yet with full gradient). The main difference between AutoAttack and vanilla PGD lies in the automatic adjustment of step size by AutoAttack, but the adjustment algorithm may be influenced by the randomness of stochastic pre-processing defense. The comparison results are presented in Tab. A2. PGD with EOT achieved significantly better attack performance across all threat models.
We additionally performed extra attack evaluation with the full-gradient C&W [33] and DeepFool [34] attacks, which are optimization-based attacks that differ significantly from PGD. Here their original $l_{2}$ setting [33, 34] were evaluated with 1,000 iteration steps and 20 EOT samples. The implementation used the standard torchattacks benchmark [^3] https://github.com/Harry24k/adversarial-attacks-pytorch. The robustness results are shown in Tab. A3 (along with the PGD and AutoAttack $l_{2}$ results in Tab. A2). We can see that PGD with EOT is also significantly stronger than other attacks such as C&W and DeepFool, which need considerably more iteration steps. Additionally, ADBM is more robust than DiffPure regardless of the evaluation methods.
@@ D.1 Comparison between Different Attack Methods on CIFAR-10
WRN-28-10 [43] AT $94.46$ $52.65$ $0.16$ $6.76$ $19.86$
WRN-28-10 [44] $93.00$ $52.70$ $0.04$ $3.27$ $18.67$
WRN-28-10 [25] $95.56$ 64.00 $0.14$ $5.07$ $23.07$
UNet+WRN-28-10 DiffPure [16] AP $93.9\pm 0.7$ $39.7\pm 2.2$ $46.1\pm 2.1$ $63.3\pm 0.8$ $49.7\pm 1.7$
@@
WRN-28-10 [43] AT $50.94$ $29.13$ $20.15$ $24.54$ $24.61$
WRN-28-10 [44] $50.89$ $27.02$ $17.79$ $23.29$ $22.70$
WRN-28-10 [25] $57.59$ 38.41 $10.73$ $27.21$ $25.45$
UNet+WRN-28-10 DiffPure [16] AP $58.0\pm 1.7$ $24.8\pm 1.8$ $44.3\pm 0.3$ $32.9\pm 1.1$ $34.0\pm 0.8$
@@
ResNet-50 [54] AT $78.8$ $47.2$ $3.6$ $25.0$ $25.3$
ResNet-50 [55] $79.5$ 50.3 $3.9$ $23.8$ $26.0$
UNet+ResNet-50 DiffPure [16] AP $77.3$ $22.2$ $55.5$ $58.4$ $45.4$
@@
D.2 Additional White-Box Results on Several Datasets
The white-box adaptive evaluation results on Tiny-ImageNet are respectively presented in Tab. A5 and Tab. A6. In the Tiny-ImageNet evaluation, we maintained consistent settings with those used for CIFAR-10, with the exception that for the $l_{\infty}$ threat, we set the bound to be $\epsilon_{\infty}=4/255$, which is a popular setting for this dataset [44, 25]. In the ImageNet-100 evaluation, we set the bounds $\epsilon_{\infty}=4/255$, $\epsilon_{1}=75$, $\epsilon_{2}=2$, which is a commonly used setting for the high-resolution datasets like ImageNet-1K. For this dataset, we directly compared ADBM with representative ResNet-50 AT checkpoints specified trained on ImageNet-1K by masking out the 900 output channels not included in ImageNet-100. By examining Tab. A5 and Tab. A6, we can draw similar conclusions to those observed on SVHN and CIFAR-10.
D.3 Additional Black-Box Results on CIFAR-10
The black-box evaluation results on CIFAR-10 are presented in Tab. A7. In this evaluation, we maintained consistent settings with those used for SVHN. By examining Tab. 4 and Tab. A7, we can conclude that under the realistic black-box attacks, ADBM has advantages over AT models even on the seen threat.
@@ D.3 Additional Black-Box Results on CIFAR-10
WRN-70-16 [41] AT $91.10$ $79.20$ $71.62$ $80.76$ $89.98$ $82.33$
WRN-70-16 [42] $88.54$ $73.91$ $69.35$ $75.82$ $87.64$ $79.05$
WRN-70-16 [24] $88.74$ $78.81$ $73.93$ $79.88$ $87.89$ $81.85$
WRN-70-16 [25] $93.25$ $81.89$ $76.30$ $82.13$ $90.34$ $84.78$
UNet+WRN-70-16 [14] AP $87.93$ $74.22$ $72.17$ $68.55$ $84.28$ $77.43$
UNet+WRN-70-16 DiffPure [16] $92.51$ $90.92$ $90.53$ $90.33$ $90.16$ $90.89$
@@
D.4 Results of ADBM on New Classifiers
Here we utilized the fine-tuned ADBM checkpoint, trained with adversarial noise from a WRN-70-16 classifier, as the pre-processor for a WRN-28-10 model and a vision transformer (ViT) model [^4] We used the checkpoint released by https://github.com/dqj5182/vit_cnn_cifar_10_from_scratch, which had 90.6% clean accuracy on CIFAR-10. directly, denoted as ADBM(Transfer). The results shown in Tab. A8 demonstrate that ADBM(Transfer) achieved robustness levels comparable to ADBM trained directly with WRN-28-10 or the ViT (52.9% vs. 53.1% on WRN-28-10, and 43.2% vs. 45.1% on ViT). This finding highlights the practicality of ADBM, as the fine-tuned ADBM model on a specific classifier can potentially be directly applied to a new classifier without requiring retraining. We guess this may be attributed to the transferability of adversarial noise [49].
@@ D.4 Results of ADBM on New Classifiers
D.5 Inference Cost of ADBM
Taking CIFAR-10 as an example, the classifier WRN-70-16 consists of 267M parameters and has an inference cost of 39 GFlops. On the other hand, the UNet architecture, specifically the DDPM++ continuous used in this work, consists of 104M parameters and has an inference cost of 18 GFlops. Therefore, although ADBM has demonstrated effectiveness even with a single reverse step (Tab. 2), the computational cost still increases by 46.2% in this case. It remains an open question whether the scale of the UNet architecture, which is currently directly adapted from architectures designed for generative tasks, can be further reduced for adversarial purification. We leave it to future work.
Appendix E Boarder Impact
Our method, which effectively purifies adversarial noise, can potentially hinder the utility of certain “adversarial for good” methods. For example, it could cause more difficult copyright protection using adversarial noise [56] when abusing our method. In addition, there seems to be a trade-off between adversarial robustness and the risk of privacy leakage [57]. But in general, we believe the concrete positive impacts on trustworthy machine learning outweigh the potential negative impacts.

### v3

@@ D.2 Additional Adaptive Attack Configurations
Appendix E Additional Evaluation Results
E.1 Comparison between Different Attack Methods on CIFAR-10
Similar to Lee & Kim (2023), we first performed a comparison between PGD with 20 EOT samples (full gradient) and AutoAttack with the same number of EOT samples (the default setting of rand version AutoAttack yet with full gradient). The main difference between AutoAttack and vanilla PGD lies in the automatic adjustment of step size by AutoAttack, but the adjustment algorithm may be influenced by the randomness of stochastic pre-processing defense. The comparison results are presented in Tab. A2. PGD with EOT achieved significantly better attack performance across all threat models.
We additionally performed extra attack evaluation with the full-gradient C&W (Carlini & Wagner, 2017) and DeepFool (Moosavi-Dezfooli et al., 2016) attacks, which are optimization-based attacks that differ significantly from PGD. Here their original $l_{2}$ setting (Carlini & Wagner, 2017; Moosavi-Dezfooli et al., 2016) were evaluated with 1,000 iteration steps and 20 EOT samples. The implementation used the standard torchattacks benchmark [^4] https://github.com/Harry24k/adversarial-attacks-pytorch. The robustness results are shown in Tab. A3 (along with the PGD and AutoAttack $l_{2}$ results in Tab. A2). We can see that PGD with EOT is also significantly stronger than other attacks such as C&W and DeepFool, which need considerably more iteration steps. Additionally, ADBM is more robust than DiffPure regardless of the evaluation methods.
@@ E.1 Comparison between Different Attack Methods on CIFAR-10
WRN-28-10 (Rade et al., 2022) AT $94.46$ $52.65$ $0.16$ $6.76$ $19.86$
WRN-28-10 (Yang et al., 2023) $93.00$ $52.70$ $0.04$ $3.27$ $18.67$
WRN-28-10 (Wang et al., 2023) $95.56$ 64.00 $0.14$ $5.07$ $23.07$
UNet+WRN-28-10 DiffPure (Nie et al., 2022) AP $93.9\pm 0.7$ $39.7\pm 2.2$ $46.1\pm 2.1$ $63.3\pm 0.8$ $49.7\pm 1.7$
@@
WRN-28-10 (Rade et al., 2022) AT $50.94$ $29.13$ $20.15$ $24.54$ $24.61$
WRN-28-10 (Yang et al., 2023) $50.89$ $27.02$ $17.79$ $23.29$ $22.70$
WRN-28-10 (Wang et al., 2023) $57.59$ 38.41 $10.73$ $27.21$ $25.45$
UNet+WRN-28-10 DiffPure (Nie et al., 2022) AP $58.0\pm 1.7$ $24.8\pm 1.8$ $44.3\pm 0.3$ $32.9\pm 1.1$ $34.0\pm 0.8$
@@
ResNet-50 (Liu et al., 2023a) AT $78.8$ $47.2$ $3.6$ $25.0$ $25.3$
ResNet-50 (Debenedetti et al., 2023) $79.5$ 50.3 $3.9$ $23.8$ $26.0$
UNet+ResNet-50 DiffPure (Nie et al., 2022) AP $77.3$ $22.2$ $55.5$ $58.4$ $45.4$
@@
WRN-70-16 (Gowal et al., 2020) AT $91.10$ $79.20$ $71.62$ $80.76$ $89.98$ $82.33$
WRN-70-16 (Rebuffi et al., 2021) $88.54$ $73.91$ $69.35$ $75.82$ $87.64$ $79.05$
WRN-70-16 (Gowal et al., 2021) $88.74$ $78.81$ $73.93$ $79.88$ $87.89$ $81.85$
WRN-70-16 (Wang et al., 2023) $93.25$ $81.89$ $76.30$ $82.13$ $90.34$ $84.78$
UNet+WRN-70-16 (Yoon et al., 2021) AP $87.93$ $74.22$ $72.17$ $68.55$ $84.28$ $77.43$
UNet+WRN-70-16 DiffPure (Nie et al., 2022) $92.51$ $90.92$ $90.53$ $90.33$ $90.16$ $90.89$
@@
E.2 Additional White-Box Results on Several Datasets
The white-box adaptive evaluation results on Tiny-ImageNet are respectively presented in Tab. A5 and Tab. A6. In the Tiny-ImageNet evaluation, we maintained consistent settings with those used for CIFAR-10, with the exception that for the $l_{\infty}$ threat, we set the bound to be $\epsilon_{\infty}=4/255$, which is a popular setting for this dataset (Yang et al., 2023; Wang et al., 2023). In the ImageNet-100 evaluation, we set the bounds $\epsilon_{\infty}=4/255$, $\epsilon_{1}=75$, $\epsilon_{2}=2$, which is a commonly used setting for the high-resolution datasets like ImageNet-1K. For this dataset, we directly compared ADBM with representative ResNet-50 AT checkpoints specified trained on ImageNet-1K by masking out the 900 output channels not included in ImageNet-100. By examining Tab. A5 and Tab. A6, we can draw similar conclusions to those observed on SVHN and CIFAR-10.
E.3 Additional Results under Black-Box Seen Threats
The evaluation results on CIFAR-10 under black-box seen threats are presented in Tab. A7. In this evaluation, we maintained consistent settings with those used for SVHN. By examining Tab. 4 and Tab. A7, we can conclude that under the realistic black-box attacks, ADBM has advantages over AT models even on the seen threat.
E.4 Results under Black-Box Unseen Threats
We further evaluate the generalization ability of ADBM on more unseen threats beyond norm-bounded attacks. Two patch-like attacks, PI-FGSM (Gao et al., 2020) and NCF (Yuan et al., 2022), and two recent diffusion-based attacks, DiffAttack (Chen et al., 2023b) and DiffPGD (Xue et al., 2023) were evaluated. As the code of these methods only natively supports input resolution $224\times 224$, we conducted experiments on ImageNet-100. The results are show in Tab. A8. We can conclude that ADBM demonstrates better generalization ability than DiffPure and previous AT methods against these black-box unseen threats.
E.5 Results of ADBM on New Classifiers
Here we utilized the fine-tuned ADBM checkpoint, trained with adversarial noise from a WRN-70-16 classifier, as the pre-processor for a WRN-28-10 model and a vision transformer (ViT) model [^5] We used the checkpoint released by https://github.com/dqj5182/vit_cnn_cifar_10_from_scratch, which had 90.6% clean accuracy on CIFAR-10. directly, denoted as ADBM (Transfer). The results shown in Tab. A9 demonstrate that ADBM (Transfer) achieved robustness levels comparable to ADBM trained directly with WRN-28-10 or the ViT (52.9% vs. 53.1% on WRN-28-10, and 43.2% vs. 45.1% on ViT). This finding highlights the practicality of ADBM, as the fine-tuned ADBM model on a specific classifier can potentially be directly applied to a new classifier without requiring retraining. We guess this may be attributed to the transferability of adversarial noise (Dong et al., 2018).
Table A8: Accuracies (%) of methods under different black-box unseen threats on ImageNet-100.
Architecture Method Type NCF PI-FGSM Diff-Attack Diff-PGD
ResNet-50 (Liu et al., 2023a) AT $3.8$ $26.3$ $21.5$ $29.4$
ResNet-50 (Debenedetti et al., 2023) $11.2$ $29.4$ $28.5$ $49.8$
UNet+ResNet-50 DiffPure AP $30.1$ $38.4$ $39.8$ $71.9$
UNet+ResNet-50 ADBM (Ours) 33.9 42.7 42.8 72.4
@@ E.5 Results of ADBM on New Classifiers
Table A10: Inference time (in seconds) for diffusion-based purification on different datasets, corresponding to the reverse sampling (RS) steps using an NVIDIA 3090 GPU (with a batch size of 1).
Dataset Network RS = 1 2 3 4 5 100 (Original DiffPure)
CIFAR-10 UNet+WRN-70-16 0.060 0.099 0.152 0.189 0.248 4.420
SVHN UNet+WRN-28-10 0.052 0.095 0.139 0.182 0.229 4.349
Tiny-ImageNet UNet+WRN-28-10 0.056 0.010 0.149 0.195 0.232 4.407
Appendix F Inference Cost of ADBM
On one hand, the inference cost of diffusion-based purification can be significantly reduced via fast sampling according to our investigation. Tab. A10 gives the actual inference time (in seconds) for diffusion-based purification on different datasets, corresponding to the reverse sampling steps. Given that ADBM has demonstrated effectiveness even with a single reverse step (Tab. 2), it has potential applications in various real-time scenarios (only 0.06s is needed for inference in this case, as shown in Tab. A10).
On the other hand, similar to DiffPure, ADBM necessitates additional parameters. Taking CIFAR-10 as an example, the classifier WRN-70-16 consists of 267M parameters, while the UNet architecture, specifically the DDPM++ continuous used in this work, consists of 104M parameters. It remains an open question whether the scale of the UNet architecture, which is currently directly adapted from architectures designed for generative tasks, can be further reduced for adversarial purification. We leave it to future work.
Appendix G Boarder Impact
Our method, which effectively purifies adversarial noise, can potentially hinder the utility of certain “adversarial for good” methods. For example, it could cause more difficult copyright protection using adversarial noise (Liang et al., 2023) when abusing our method. In addition, there seems to be a trade-off between adversarial robustness and the risk of privacy leakage (Li et al., 2024a). But in general, we believe the concrete positive impacts on trustworthy machine learning outweigh the potential negative impacts.